OpenStudy (anonymous):

Add negative seven-eighths plus one-fourth. negative seven thirty-seconds negative one and one-eighth negative five-eighths negative three and one-half

OpenStudy (shane_b):

Do you know where to start?

OpenStudy (anonymous):

do i need to find the mcd?

OpenStudy (shane_b):

Well, I call it the lcd (lowest common denominator) but yes, you need to find that.

OpenStudy (anonymous):

and 8 can be the ICD right?

OpenStudy (shane_b):

Yes

OpenStudy (shane_b):

Hint:\[(\frac{1}{4})(\frac{2}{2})=\frac{2}{8}\]

OpenStudy (anonymous):

well i did it and i got 9/8 and i need to make it a proper fraction

OpenStudy (shane_b):

You didn't account for the fact that it's NEGATIVE 7/8

OpenStudy (shane_b):

\[\large \frac{-7}{8}+\frac{2}{8}=?\]

OpenStudy (shane_b):

So -7 + 2 = ?

OpenStudy (anonymous):

so -5/8

OpenStudy (shane_b):

Yep :)

OpenStudy (anonymous):

thank you so much

OpenStudy (shane_b):

no problem, good luck!